# Millimeter-wave Radio SLAM: End-to-End Processing Methods and Experimental Validation

### v1

@@ III-A BRSRP Measurements
[…]
$\mathbf{b}_{n_{1}}^{\mathrm{H}}\mathbf{b}_{n_{2}}\approx 0~{}~{}~{}{\rm{or}}~{}~{}~{}\mathbf{c}_{n_{1}}^{\mathrm{H}}\mathbf{c}_{n_{2}}\approx 0$ (5)
@@
$\beta_{i,j}\approx\!\sum_{n=1}^{N}|\xi_{n}|^{2}|G_{\mathrm{TX},i}(\boldsymbol{\psi}_{\mathrm{TX},n})|^{2}|G_{\mathrm{RX},j}(\boldsymbol{\psi}_{\mathrm{RX},n})|^{2}\!+\!\bar{N}_{i,j}$ (6)
where $\bar{N}_{i,j}=(1/N_{\mathrm{RS}})\sum_{(k,m)\in\mathcal{M}_{\mathrm{RS}}}|\tilde{n}_{k,m}^{i,j}|^{2}$ denotes the averaged noise power.
@@ III-B1 Technical Premise and Intuition
where $\mathbf{g}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})$ and $\mathbf{g}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})$ denote the TX and RX beam gain vectors, with $\mathbf{g}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})[i]=|G_{\text{TX},i}(\boldsymbol{\psi}_{\text{TX},n})|^{2}$ and $\mathbf{g}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})[j]=|G_{\text{RX},j}(\boldsymbol{\psi}_{\text{RX},n})|^{2}$, respectively. Furthermore, $\bar{\mathbf{N}}$ is a noise matrix whose element at the $i^{\text{th}}$ row and $j^{\text{th}}$ column is defined as $\bar{\mathbf{N}}[i,j]=\bar{N}_{i,j}$.
@@
Inspired by the structure of $\mathbf{B}$ in (8), let us consider its singular value decomposition (SVD), expressed as $\mathbf{B}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{L_{\textrm{TX}}\times L_{\textrm{TX}}}$, $\mathbf{V}\in\mathbb{R}^{L_{\textrm{RX}}\times L_{\textrm{RX}}}$ are unitary matrices, and $\mathbf{\Lambda}\in\mathbb{R}^{L_{\textrm{TX}}\times L_{\textrm{RX}}}$ is a rectangular diagonal matrix. In the main diagonal of $\mathbf{\Lambda}$, singular values of $\mathbf{B}$, $[\sigma_{1},\sigma_{2},...,,\sigma_{R}]>0$ are sorted in the descending order, where $R=\text{rank}(\mathbf{B})\leq\text{min}(L_{\textrm{TX}},L_{\textrm{RX}})$. The SVD can also be expressed as
@@
where $q_{r,n}^{\text{u}}$ and $q_{r,n}^{\text{v}}$ are the linear combination weights of the spanning set vectors $\mathbf{g}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})$ and $\mathbf{g}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})$, respectively. In addition, $\mathbf{B}^{(1)}_{r}$ in (9) is a rank-1 matrix associated with the $r^{\text{th}}$ singular value and the related basis vectors $\mathbf{u}_{r}$ and $\mathbf{v}_{r}$. Furthermore, in the case of a sparse beamformed channel where all paths are well separated, the following holds
@@
Interestingly, based on (8)–(10) and the corresponding fact that the singular vectors are in the span of the beam gain vectors, the SVD-based low-rank modeling approach can also inherently manage different unknown beam patterns, and is thus resilient to sidelobes and other beam pattern fluctuations. Specifically, the antenna sidelobes illustrated also in Fig. 3, can be decoupled from the peaks in $\mathbf{B}$ by using different low-rank approximations with increasing rank and analyzing the resulting power increment formed at each step. To this end, a pair of subsequent low-rank approximations of $\mathbf{B}$, denoted as $\mathbf{B}_{K-1}$ and $\mathbf{B}_{K}$, with $K<R$, can be expressed as
@@ III-B2 Overall Approach and Refinements
Input: TX beam angles $\{\Phi_{i}\}_{i=1}^{L_{\text{TX}}}$, RX beam angles $\{\Theta_{j}\}_{j=1}^{L_{\text{RX}}}$, power measurement matrix $\mathbf{B}\in\mathbb{R}^{L_{\text{TX}}\times L_{\text{RX}}}$, desired power ratio $p$, and power threshold $\beta_{th}$
@@
12:     Compute the power ratio: $p_{r}\leftarrow\frac{1}{p_{\text{TOT}}}\sum_{\tilde{r}=1}^{r}\sigma_{\tilde{r}}^{2}$
@@
16:         $\mathcal{M}_{\text{est}}\leftarrow\{\mathcal{M}_{\text{est}}\,|\,\text{elements with }\beta_{\hat{i}_{r},\hat{j}_{r}}\geq\beta_{th}\}$
@@
The complete proposed procedure for obtaining AoD estimates $\{\hat{\phi}_{n}\}_{n=1}^{\hat{N}}$ and AoA estimates $\{\hat{\theta}_{n}\}_{n=1}^{\hat{N}}$ for $\hat{N}$ estimated paths, is stated and described in Algorithm 1. Besides the fundamental processing of rank-1 matrices for extracting angle information, three additional refinements are shown and applied, namely, power thresholding, clustering, and polynomial fitting. With power thresholding, we omit peaks which are close to noise level, and thus mitigate noise-related estimation errors. The clustering step, in turn, considers the fact that powerful peaks create large residuals and require more than one rank-1 singular matrix to be properly represented. Assuming spatially sparse channel, we expect such residual peaks to occur only in vicinity of powerful peaks that are spatially separated from each other, thus creating clusters in AoD-AoA domain. After the clustering step, where any existing clustering method [39] can be adopted, $\hat{N}$ clusters are obtained, whose cluster means represent the coarsely estimated AoD and AoA angles. Lastly, to enhance the angle estimation precision, each peak associated with a coarsely estimated AoD and AoA is fitted with a parabolic surface using a weighted least squares method. Fitting is performed within a local window of size approximately representing the beamwidth. The final AoD estimates $\{\hat{\phi}_{n}\}_{n=1}^{\hat{N}}$ and AoA estimates $\{\hat{\theta}_{n}\}_{n=1}^{\hat{N}}$ can be then found in closed form as a vertex of the parabolic surface for each peak.
@@ III-B3 Complexity Assessment
The fundamental complexity order of the proposed SVD-based angle estimation method is $\mathcal{O}(\text{min}(L_{\textrm{TX}}L_{\textrm{RX}}^{2},L_{\textrm{TX}}^{2}L_{\textrm{RX}})+\text{min}(L_{\textrm{TX}},L_{\textrm{RX}})^{2}K+L_{\textrm{TX}}L_{\textrm{RX}}K)$, excluding further refinements of thresholding, clustering and polynomial fitting. The corresponding complexities of [15] as well as a classical cell-averaging constant false alarm rate (CFAR) detector [40], being used as benchmark methods along the numerical results, read $\mathcal{O}(\text{min}(L_{\textrm{TX}},L_{\textrm{RX}})^{3}+N_{\text{peak}}L_{\textrm{TX}}L_{\textrm{RX}})$, and $\mathcal{O}(L_{\textrm{TX}}L_{\textrm{RX}}N_{\text{TB}}^{2})$, respectively, where $N_{\text{peak}}$ is a maximum number of support squares and $N_{\text{TB}}$ is a size of the CFAR training window. For good performance, $N_{\text{TB}}^{2}$ is comparable to or larger than $L$ (i.e., $L_{\textrm{TX}}$ or $L_{\textrm{RX}}$), and thus all three methods have similar complexity order of $\mathcal{O}(L^{3})$.
@@ III-C1 Coarse ToA Estimation in Time Domain
Furthermore, the coarse delay estimate, defined with respect to the TX reference time, reads then $\hat{\tau}^{i,j}_{\text{c}}=\widehat{\Delta q}_{i,j}/F_{s}$.
@@ III-C2 Fine ToA Estimation in Frequency Domain
The fine time delay estimation is performed per estimated path, using the frequency domain samples $y_{k,m}^{i,j}$ with $(k,m)\in\mathcal{M}_{RS}$, corresponding to the beam pair $(i,j)$ whose beam angles are closest to the estimated AoD and AoA. Such fine ToA estimate for path $n$, determined with respect to the beginning of the OFDM symbol, can be obtained as [42, Ch. 3.2]
@@
Finally, the complete ToA estimates are obtained as
$\hat{\tau}_{n}=\hat{\tau}^{\hat{i}_{n},\hat{j}_{n}}_{\text{c}}+\hat{\tau}_{\text{f},n}$ (15)
for path indices $n=1,\dots,\hat{N}$. As noted in [43, 41], the coarse ToA estimate can also be determined wrt. the frame start time. In such case, the possible excess time between the beginning of the frame and the actual RS transmission time can be taken into account in the coarse ToA estimate, while the path-wise fine ToA estimates remains intact. With unsynchronized TX and RX clocks, the ToA estimate in (15) is affected by the clock bias, as considered in Section IV.
@@ IV-A Problem Formulation
It is assumed that the BS position and orientation, denoted with $[{\mathbf{p}}_{\textrm{BS}}^{\top},\,\alpha_{\textrm{BS}}]^{\top}$, are known while the unknown UE state is represented using the 2D position, heading and clock bias (cast in meters) as ${\mathbf{s}}=[{\mathbf{p}}_{\textrm{UE}}^{\top},\,\alpha_{\textrm{UE}},\,B_{\textrm{UE}}]^{\top}$. Furthermore, the $n$th single bounce propagation path is represented using the 2D interaction point or landmark ${\mathbf{m}}_{n}=[x_{n},\,y_{n}]^{\top}$, $n=2,3,...,\hat{N}$, where $\hat{N}$ denotes the number of the available AoA, AoD and ToA estimates in a given measurement location. For notational convenience, the LoS path index – if existing – is $n=1$.
@@
where $p({\mathbf{x}})=\mathcal{N}({\mathbf{x}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$ is the prior for ${\mathbf{x}}$ obtained for example using external sensors or a previous estimate and $p({\mathbf{z}}_{n}\mid{\mathbf{x}}_{n})$ is the likelihood of the $n$th measurement. It is to be noted that typically snapshot SLAM implies that no prior information exists since the primary interest is analyzing what can be done with radio signals alone [16, 28] whereas in this article, we evaluate the system performance both with and without the prior. The measurements are defined as ${\mathbf{z}}_{n}=[c\hat{\tau}_{n},\,\hat{\phi}_{n},\,\hat{\theta}_{n}]^{\top}$ in which the delay estimates are converted to meters for convenience. An estimate for ${\mathbf{x}}$ can be obtained by maximizing (16), mathematically given by
@@ IV-B Measurement Models
${\mathbf{h}}_{n}({\mathbf{x}})=\begin{bmatrix}d+B_{\textrm{UE}}\\ \textrm{atan}2(-\delta_{1,y},-\delta_{1,x})-\alpha_{\textrm{BS}}\\ \textrm{atan}2(\delta_{2,y},\delta_{2,x})-\alpha_{\textrm{UE}}\end{bmatrix}.$ (19)
For the LoS path $(n=1)$, the parameters are defined as: $d=\lVert{\mathbf{p}}_{\textrm{BS}}-{\mathbf{p}}_{\textrm{UE}}\rVert$ and $[\delta_{1,x},\,\delta_{1,y}]^{\top}=[\delta_{2,x},\,\delta_{2,y}]^{\top}={\mathbf{p}}_{\textrm{BS}}-{\mathbf{p}}_{\textrm{UE}}$. Respectively for the $n$th NLoS path, the parameters are defined as: $d=\lVert{\mathbf{p}}_{\textrm{BS}}-{\mathbf{m}}_{n}\rVert+\lVert{\mathbf{m}}_{n}-{\mathbf{p}}_{\textrm{UE}}\rVert$, $[\delta_{1,x},\,\delta_{1,y}]^{\top}={\mathbf{p}}_{\textrm{BS}}-{\mathbf{m}}_{n}$ and $[\delta_{2,x},\,\delta_{2,y}]^{\top}={\mathbf{m}}_{n}-{\mathbf{p}}_{\textrm{UE}}$.

### v2

@@ III-A BRSRP Measurements
Let $\mathbf{b}_{n}\in\mathbb{C}^{K}$ and $\mathbf{c}_{n}\in\mathbb{C}^{M}$ denote the frequency-domain and time-domain steering vectors of the $n^{\text{th}}$ path, respectively, where $K$ is the number of active RS subcarriers, and $M$ is the number of RS OFDM symbols. Moreover, the $k^{\text{th}}$ element of $\mathbf{b}_{n}$ is defined as $\mathbf{b}_{n}[k]=e^{-\mathrm{i}2\pi k\Delta f\tau_{\mathrm{f},n}}$, and the $m^{\text{th}}$ element of $\mathbf{c}_{n}$ as $\mathbf{c}_{n}[m]=e^{\mathrm{i}2\pi mT_{\text{sym}}f_{\text{D},n}}$. Suppose that the paths in (2) are non-overlapping in either delay or Doppler, and that $K$ and/or $M$ are/is sufficiently large, i.e.,
$\frac{\mathbf{b}_{n_{1}}^{\mathrm{H}}\mathbf{b}_{n_{2}}}{{K}}\approx 0~{}~{}~{}{\rm{or}}~{}~{}~{}\frac{\mathbf{c}_{n_{1}}^{\mathrm{H}}\mathbf{c}_{n_{2}}}{{M}}\approx 0$ (6)
@@
$\!\!\beta_{i,j}\!\approx\!\sum_{n=1}^{N}\!|\xi_{n}|^{2}|G_{\mathrm{TX},i}(\boldsymbol{\psi}_{\mathrm{TX},n})|^{2}|G_{\mathrm{RX},j}(\boldsymbol{\psi}_{\mathrm{RX},n})|^{2}+{\sigma_{\mathrm{noise}}^{2}}$ (7)
where $\sigma_{\mathrm{noise}}^{2}$ is the variance of $\tilde{n}_{k,m}^{i,j}$.
@@ III-B1 Technical Premise and Intuition
where $\mathbf{g}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})$ and $\mathbf{g}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})$ denote the TX and RX beam gain vectors, with $\mathbf{g}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})[i]=|G_{\text{TX},i}(\boldsymbol{\psi}_{\text{TX},n})|^{2}$ and $\mathbf{g}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})[j]=|G_{\text{RX},j}(\boldsymbol{\psi}_{\text{RX},n})|^{2}$, respectively. Furthermore, $\bar{\mathbf{N}}$ is a noise matrix whose element at the $i^{\text{th}}$ row and $j^{\text{th}}$ column is defined as $\bar{\mathbf{N}}[i,j]={(1/N_{\mathrm{RS}})\!\sum_{(k,m)\in\mathcal{M}_{\mathrm{RS}}}\!|\tilde{n}_{k,m}^{i,j}|^{2}}$.
@@
Inspired by the structure of $\mathbf{B}$ in (9), let us consider its singular value decomposition (SVD), expressed as $\mathbf{B}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{L_{\textrm{TX}}\times L_{\textrm{TX}}}$, $\mathbf{V}\in\mathbb{R}^{L_{\textrm{RX}}\times L_{\textrm{RX}}}$ are unitary matrices, and $\mathbf{\Lambda}\in\mathbb{R}^{L_{\textrm{TX}}\times L_{\textrm{RX}}}$ is a rectangular diagonal matrix. In the main diagonal of $\mathbf{\Lambda}$, singular values of $\mathbf{B}$, $[\sigma_{1},\sigma_{2},...,,\sigma_{R}]>0$, are sorted in the descending order, where $R=\text{rank}(\mathbf{B})\leq\text{min}(L_{\textrm{TX}},L_{\textrm{RX}})$. The SVD can also be expressed as
@@
where $q_{r,n}^{\text{u}}$ and $q_{r,n}^{\text{v}}$ are the linear combination weights of the spanning set vectors $\mathbf{g}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})$ and $\mathbf{g}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})$, respectively. In addition, $\mathbf{B}^{(1)}_{r}$ in (10) is a rank-1 matrix associated with the $r^{\text{th}}$ singular value and the related basis vectors $\mathbf{u}_{r}$ and $\mathbf{v}_{r}$. Furthermore, in the case of a sparse beamformed channel where all paths are well separated, combined with directional beamforming at the TX and RX, the following holds [40, 37]
@@
Interestingly, based on (9)–(11) and the corresponding fact that the singular vectors of the BRSRP matrix are in the span of the beam gain vectors, the SVD-based low-rank modeling approach can also inherently manage different unknown beam patterns, and is thus resilient to sidelobes and other beam pattern fluctuations. Specifically, the antenna sidelobes illustrated also in Fig. 3, can be decoupled from the peaks in $\mathbf{B}$ by using different low-rank approximations with increasing rank and analyzing the resulting power increment formed at each step. To this end, a pair of subsequent low-rank approximations of $\mathbf{B}$, denoted as $\mathbf{B}_{K-1}$ and $\mathbf{B}_{K}$, with $K<R$, can be expressed as
@@ III-B2 Overall Approach and Refinements
Input: TX beam angles $\{\Phi_{i}\}_{i=1}^{L_{\text{TX}}}$, RX beam angles $\{\Theta_{j}\}_{j=1}^{L_{\text{RX}}}$, power measurement matrix $\mathbf{B}\in\mathbb{R}^{L_{\text{TX}}\times L_{\text{RX}}}$, desired power ratio $p$, and power threshold $\beta_{\text{th}}$
@@
12:    Compute power ratio: $p_{r}\leftarrow\frac{1}{p_{\text{TOT}}}\sum_{\tilde{r}=1}^{r}\sigma_{\tilde{r}}^{2}\times 100\%$
@@
16:         $\mathcal{M}_{\text{est}}\leftarrow\{\mathcal{M}_{\text{est}}\,|\,\text{elements with }\beta_{\hat{i}_{r},\hat{j}_{r}}\geq\beta_{\text{th}}\}$
@@
The complete proposed procedure for obtaining AoD estimates $\{\hat{\phi}_{n}\}_{n=1}^{\hat{N}}$ and AoA estimates $\{\hat{\theta}_{n}\}_{n=1}^{\hat{N}}$ for $\hat{N}$ estimated paths, is stated and described in Algorithm 1. Besides the fundamental processing of rank-1 matrices for extracting angle information, three additional refinements are shown and applied, namely, power thresholding, clustering, and polynomial fitting. With power thresholding, we omit peaks which are close to noise level, and thus mitigate noise-related estimation errors. One feasible approach to assign a value for the related power threshold $\beta_{\text{th}}$ is to set it at or marginally higher than the prevailing noise floor. Concrete examples are provided along the numerical results, in Section VI. The clustering step, in turn, considers the fact that powerful peaks create large residuals and require more than one rank-1 singular matrix to be properly represented. Assuming spatially sparse channel, we expect such residual peaks to occur only in vicinity of powerful peaks that are spatially separated from each other, thus creating clusters in AoD-AoA domain. After the clustering step, where any existing clustering method [41] can be adopted, $\hat{N}$ clusters are obtained, whose cluster means represent the coarsely estimated AoD and AoA angles. Lastly, to enhance the angle estimation precision, each peak associated with a coarsely estimated AoD and AoA is fitted with a parabolic surface using a weighted least squares method. Fitting is performed within a local window of size approximately representing the beamwidth. The final AoD estimates $\{\hat{\phi}_{n}\}_{n=1}^{\hat{N}}$ and AoA estimates $\{\hat{\theta}_{n}\}_{n=1}^{\hat{N}}$ can be then found in closed form as a vertex of the parabolic surface for each peak.
@@ III-B3 Complexity Assessment
The fundamental complexity order of the proposed SVD-based angle estimation method is $\mathcal{O}(\text{min}(L_{\textrm{TX}}L_{\textrm{RX}}^{2},L_{\textrm{TX}}^{2}L_{\textrm{RX}})+\text{min}(L_{\textrm{TX}},L_{\textrm{RX}})^{2}K+L_{\textrm{TX}}L_{\textrm{RX}}K)$, excluding further refinements of thresholding, clustering and polynomial fitting. The corresponding complexities of the method in [15] as well as that of a classical cell-averaging constant false alarm rate (CFAR) detector [42], being used as benchmark methods along the numerical results, read $\mathcal{O}(\text{min}(L_{\textrm{TX}},L_{\textrm{RX}})^{3}+N_{\text{peak}}L_{\textrm{TX}}L_{\textrm{RX}})$, and $\mathcal{O}(L_{\textrm{TX}}L_{\textrm{RX}}N_{\text{TB}}^{2})$, respectively, where $N_{\text{peak}}$ is a maximum number of support squares [15] and $N_{\text{TB}}$ is a size of the CFAR training window. For good performance, $N_{\text{TB}}^{2}$ is comparable to or larger than $L$ (i.e., $L_{\textrm{TX}}$ or $L_{\textrm{RX}}$), and thus all three methods have similar complexity order of $\mathcal{O}(L^{3})$.
@@ III-C1 Coarse ToA Estimation in Time Domain
Furthermore, the coarse delay estimate reads then
$\hat{\tau}^{i,j}_{\text{c}}=\widehat{\Delta q}_{i,j}/F_{s},$ (15)
which defines the delay between the received OFDM symbol and the transmit time with respect to the receiver clock – and is thus subject to bias.
@@ III-C2 Fine ToA Estimation in Frequency Domain
The fine time delay estimation is performed per estimated path, using the frequency domain samples $y_{k,m}^{i,j}$ with $(k,m)\in\mathcal{M}_{RS}$, corresponding to the beam pair $(i,j)$ whose beam angles are closest to the estimated AoD and AoA. Such fine ToA estimate for path $n$, determined with respect to the beginning of the OFDM symbol, can be obtained as [44, Ch. 3.2]
@@
Finally, the complete (biased) ToA estimates are obtained as
${\hat{\tau}^{b}_{n}}=\hat{\tau}^{\hat{i}_{n},\hat{j}_{n}}_{\text{c}}+\hat{\tau}_{\text{f},n}$ (17)
for path indices $n=1,\dots,\hat{N}$, which correspond to (biased) path distance estimates of $c\hat{\tau}^{b}_{n}$. As noted in [45, 43], the coarse ToA estimate can also be determined wrt. the frame start time. In such case, the possible excess time between the beginning of the frame and the actual RS transmission time can be taken into account in the coarse ToA estimate, while the path-wise fine ToA estimates remains intact. With unsynchronized TX and RX clocks, the underlying bias of ToA estimate in (17) can be estimated and dealt with in the considered SLAM scenario, as shown in Section IV.
@@ IV-A Problem Formulation
It is assumed that the BS position and orientation, denoted with $[{\mathbf{p}}_{\textrm{BS}}^{\top},\,\alpha_{\textrm{BS}}]^{\top}$, are known while the unknown UE state is represented using the 2D position, heading and clock bias (cast in meters, $B_{\textrm{UE}}=c\hskip 0.85358ptb_{\textrm{UE}}$) as ${\mathbf{s}}=[{\mathbf{p}}_{\textrm{UE}}^{\top},\,\alpha_{\textrm{UE}},\,B_{\textrm{UE}}]^{\top}$. Furthermore, the $n$th single bounce propagation path is represented using the 2D interaction point or landmark ${\mathbf{m}}_{n}=[x_{n},\,y_{n}]^{\top}$, $n=2,3,...,\hat{N}$, where $\hat{N}$ denotes the number of the available AoA, AoD and ToA estimates in a given measurement location. For notational convenience, the LoS path index – if existing – is $n=1$.
@@
where $p({\mathbf{x}})=\mathcal{N}({\mathbf{x}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$ is the prior for ${\mathbf{x}}$ obtained for example using external sensors or a previous estimate and $p({\mathbf{z}}_{n}\mid{\mathbf{x}}_{n})$ is the likelihood of the $n$th measurement. It is to be noted that typically snapshot SLAM implies that no prior information exists since the primary interest is analyzing what can be done with radio signals alone [16, 28] whereas in this article, we evaluate the system performance both with and without the UE prior. No prior knowledge on the map is required, but for a comprehensive treatment of the problem, the solution is presented considering both the UE and map priors. The measurements are defined as ${\mathbf{z}}_{n}=[{c\hat{\tau}^{b}_{n}},\,\hat{\phi}_{n},\,\hat{\theta}_{n}]^{\top}$, in which the delay estimates are converted to meters for convenience. An estimate for ${\mathbf{x}}$ can be obtained by maximizing (18), mathematically given by
@@ IV-B Measurement Models
${\mathbf{h}}_{n}({\mathbf{x}})=\begin{bmatrix}d{-}B_{\textrm{UE}}\\ \textrm{atan}2(-\delta_{1,y},-\delta_{1,x})-\alpha_{\textrm{BS}}\\ \textrm{atan}2(\delta_{2,y},\delta_{2,x})-\alpha_{\textrm{UE}}\end{bmatrix}.$ (21)
For the LoS path $(n=1)$, the parameters are defined as: $d=\lVert{\mathbf{p}}_{\textrm{BS}}-{\mathbf{p}}_{\textrm{UE}}\rVert$ and $[\delta_{1,x},\,\delta_{1,y}]^{\top}=[\delta_{2,x},\,\delta_{2,y}]^{\top}={\mathbf{p}}_{\textrm{BS}}-{\mathbf{p}}_{\textrm{UE}}$. Respectively for the $n$th NLoS path, the parameters are defined as: $d=\lVert{\mathbf{p}}_{\textrm{BS}}-{\mathbf{m}}_{n}\rVert+\lVert{\mathbf{m}}_{n}-{\mathbf{p}}_{\textrm{UE}}\rVert$, $[\delta_{1,x},\,\delta_{1,y}]^{\top}={\mathbf{p}}_{\textrm{BS}}-{\mathbf{m}}_{n}$, and $[\delta_{2,x},\,\delta_{2,y}]^{\top}={\mathbf{m}}_{n}-{\mathbf{p}}_{\textrm{UE}}$.